OpenStudy (anonymous):

Solve the equation. Write all proposed solutions. If a solution is extraneous, so indicate. square root of b^2 -b = square root of 3-b^2

OpenStudy (freckles):

\[\sqrt{b^2-b}=\sqrt{3-b^2} \\ \text{ solve } b^2-b=3-b^2 \]

OpenStudy (freckles):

It is a quadratic if you didn't notice. Write in the form ax^2+bx+c=0 and use the quadratic formula or you can try to factor

OpenStudy (anonymous):

okay thanks

OpenStudy (freckles):

after you find b check both solutions